(1 points) Question 6 A single-turn circular loop of wire of radius 5.0 cm lies in a plane perpendicular to a spatially unifoPLEASE SHOW WORK! :)
Answer #1

Flux Houqu a lоор и 9 vеи by ф» С.А . - ф•лэ? Emf i ... E= -do i Lenz low). аф = Ф- 9) ф, - В, А ф = $, А . do= A (B2B) - п